# Multi-Scale Contrastive Learning for Video Temporal Grounding

### v2

@@ Introduction
Temporal video grounding aims to localize moments of interest in an untrimmed video given a free-form textual description. It is a challenging multimodal task since it involves understanding temporal information in videos and reasoning about their connections to semantic information in texts. Recently, temporal grounding has drawn increasing attention (Mu, Mo, and Li 2024; Jung et al. 2023; Xu et al. 2023; Pan et al. 2023), due to its wide range of applications such as surveillance (Zhang, Zhu, and Roy-Chowdhury 2016), robotics (Burgner-Kahrs, Rucker, and Choset 2015), and autonomous driving (Claussmann et al. 2019).
@@
Previous methods (Zhang et al. 2020b; Soldan et al. 2021; Zhang et al. 2020a) for temporal grounding concentrate on grounding merely a few queries in short video snippets. However, recently the growing availability of long videos, e.g. on streaming platforms, and demands to query their rich content have necessitated productive grounding of large volumes of queries in long videos. Because of such short-to-long video paradigm shift, latest methods (Zhang, Wu, and Li 2022; Mu, Mo, and Li 2024) have utilized local self-attention to restrict attention within a local window, following the intuition that temporal context beyond a certain range is less helpful for moment localization.
@@
To capture moments at different temporal scales without enlarging the window size of the local self-attention, recent methods (Zhang, Wu, and Li 2022; Mu, Mo, and Li 2024) need to combine several Transformer blocks with downsampling between every two blocks, resulting in a feature pyramid of moment representations, as illustrated in Figure 1 (left). Unfortunately, due to such downsampling operation, when moment representations are propagated from lower levels of short-range (local) moments to higher levels of long-range (global) moments, information contained in representations of longer moments will gradually degrade (Guo et al. 2020; Yang et al. 2023). This could explain why performance of these methods tends to degrade as the duration of target moments increase, as shown in Figure 1 (right) and statistically shown with Intersection-over-Union (IoU) results in Figure 2, respectively.
@@
Figure 1: (Left) Illustration of feature pyramid to encode video moments of different lengths; (Right) An example where recent method SnAG (Mu, Mo, and Li 2024) accurately localizes short video moment but fails on long moment.
@@
Figure 2: First and Second: IoU results with respect to target video moment length on Ego4D-NLQ (Grauman et al. 2022) of baseline SnAG (Mu, Mo, and Li 2024) and our model. Third and Fourth: IoU results with respect to target video moment length on TACoS (Regneri et al. 2013) datasets of baseline SnAG (Mu, Mo, and Li 2024) and our model.
@@
To enrich information in video moment representations, recent works (Panta et al. 2024; Xiao et al. 2024; Ji et al. 2024; Liu et al. 2024) have employed contrastive learning for temporal grounding. The intuition is to capture mutual information between video moments and textual query to preserve salient semantics in moment representations. These works mainly involve query-moment pairs in which queries relate to video moments of distinct videos, hence the learned semantics among moment representations would be independent from each other. However, such approach might not be suitable for the latest scalable video-centric approach (Zhang, Wu, and Li 2022; Mu, Mo, and Li 2024), in which multiple textual queries are related to one video. Therefore, if the grounding of two textual queries results in temporal overlapping, there might be a conflict in compact moment representations (An et al. 2023). Furthermore, focusing upon moment-query relations limits these works to the feature space of the final encoder layer, which could not effectively utilize all hidden representations across encoding layers. For multi-scale temporal grounding, such cross-scale representations should be fully used since they express semantics in video moments of various lengths.
@@
Temporal Grounding. Research works on temporal grounding can be categorized into two groups: two-stage and single-stage. In the two-stage group, methods generate temporal segments as proposals, then score the segments’ probabilities of being target moments and predict the refined boundary timestamps. Early approaches (Anne Hendricks et al. 2017; Gao et al. 2017) densely sample proposals leveraging sliding windows and score the proposals independently. Instead of independent sampling, Liu et al. (2021); Xiao et al. (2021a, b) subsequently condition proposal generation upon sentence queries and/or video context to avoid dense sampling. In contrast, Gao et al. (2021); Soldan et al. (2021); Wang et al. (2021a); Zhang et al. (2021) enumerate all segments and organize them into a 2D adjacency map for relation prediction. In the single-stage group, methods localize moments in a single shot without utilizing proposals, thus being more efficient than the two-stage group. Several works decode moment boundaries from a pooled representation (Li et al. 2022; Li, Guo, and Wang 2021; Zhou et al. 2021) or learnable queries (Nguyen et al. 2023b; Lin et al. 2023; Woo et al. 2022). Most related to our work are models that denote moment candidates as points (Fang et al. 2023; Liu et al. 2022; Mu, Mo, and Li 2024; Li, Guo, and Wang 2021). To capture temporal moments of various lengths, these methods designate a feature pyramid to produce moment representations of various scales, resulting in semantic gaps for longer-length moments.
Contrastive Learning. Notable improvements contributed by contrastive learning are made by contrastive loss applied to the final encoded outputs (Nguyen and Luu 2021; Wu et al. 2023; Nguyen et al. 2025; Hu, Cui, and Wang 2021; Nguyen et al. 2022; Wu et al. 2024; Nguyen et al. 2024b; Wang et al. 2021b; Nguyen et al. 2023a, 2024a). Wang et al. (2021b); Hu, Cui, and Wang (2021) employ a memory bank to maintain an extended set of positive and negative samples. Instead of utilizing outputs at the single final layer, contrastive learning with local and global representations across different layers has been widely studied (Hjelm et al. 2018; Zhang et al. 2020c; Bachman, Hjelm, and Buchwalter 2019; Chaitanya et al. 2020). Zhang et al. (2020c) maximize the mutual information between representations of different local windows of a sentence and the representation of the global sentence. In addition, Bachman, Hjelm, and Buchwalter (2019) propose an InfoNCE loss which optimizes over local and global features of two augmented views of an image, whereas Chaitanya et al. (2020) specifically concentrate on medical images.
@@
$\bar{Z}^{l}=\alpha^{l}\cdot\text{LocalMSA}\left(\text{LN}\left(Z^{l-1}\right)\right)+Z^{l-1},$ (2)
@@
Inspired by (Mu, Mo, and Li 2024), we implement the downsampling operator $\downarrow$ as a strided depthwise 1D convolution. The downsampling operation engenders the multi-scale property of the encoder, generating representations for longer video moments.
@@
where $\mathcal{U}$ denotes a discrete uniform distribution, $\{y_{j^{\prime}}^{l}\}_{l=1}^{L}$ the set of target video moments in each layer $l$, and $N_{Q}$ the number of textual queries related to video $V$. We generate the target set $\mathcal{P}(l)$ via center sampling (Zhang, Wu, and Li 2022; Mu, Mo, and Li 2024), i.e. given any moment centered at $t$, any time step $c\in\left[t-\alpha\frac{T}{T^{l}},t+\alpha\frac{T}{T^{l}}\alpha\right]$ in layer $l$ is considered as a target. After sampling the query and target moments, we directly utilize the representations $\{\mathbf{z}_{j^{\prime}}^{l}\}_{l=1}^{L}$ of the target moments $\{y_{j^{\prime}}^{l}\}_{l=1}^{L}$ extracted by the aforementioned multi-scale video encoder.
@@
To fairly compare with previous works and satisfy the scalability of temporal grounding operation for long videos, we adopt video-centric sampling approach (Mu, Mo, and Li 2024). For Ego4D-NLQ, we use pre-trained 1) SlowFast video features (Feichtenhofer et al. 2019) with BERT textual features (Devlin et al. 2018), and 2) EgoVLP video and textual features (Lin et al. 2022). For testing, we report R@$\{1,5\}$, tIoU = $\{0.3,0.5\}$. For MAD dataset, we use CLIP features (Radford et al. 2021) for both videos and texts, and report R@$\{1,5,10,50\}$, tIoU = $\{0.1,0.3,0.5\}$. For the TACoS dataset, we use C3D video features (Tran et al. 2015) and GloVe textual features (Pennington, Socher, and Manning 2014). We report results in terms of R@$\{1,5\}$, tIoU = $\{0.5,0.7\}$. In addition, we utilize I3D features (Carreira and Zisserman 2017) pre-trained on Kinetics (Kay et al. 2017) for Charades-STA and C3D features (Tran et al. 2015) for ActivityNet-Captions experiments. For both datasets, similar to TACoS, we take advantage of GloVe textual features (Pennington, Socher, and Manning 2014). We report R@$\{1,5\}$, tIoU = $\{0.5,0.7\}$ for testing on Charades-STA, and R@$\{1,5\}$, tIoU = $\{0.3,0.5\}$ for testing on ActivityNet-Captions. For more details regarding model architecture, we direct interested readers to the appendix. For both within-scale and cross-scale contrastive learning implementation, we keep the size of the negative sample set $\mathcal{N}(l)$ in every level $l$ to be equal to the size of the positive video clips $\mathcal{P}(l)$ that correspond to the target video moments. Based upon validation and fair comparison with previous methods, we use $\rho_{\text{ref}}=\rho_{\text{within}}=\rho_{\text{cross}}$ = 1.0.
@@
We consider the following temporal grounding models as baselines: (i) VSL-Net (Zhang et al. 2020a) utilizing textual query to highlight regions potential to comprise the target moment; (ii) VLG-Net (Soldan et al. 2021) modeling temporal grounding as a graph matching problem; (iii) Moment-DETR (Lei, Berg, and Bansal 2021), a Transformer encoder-decoder architecture that views temporal grounding as a set prediction problem; (iv) CONE (Hou et al. 2022) subsequently slicing a video input into windows, selects relevant windows, and ranks the selected windows to obtain target moments; (v) MMN (Wang et al. 2022), a Siamese-like network architecture that is trained with video-query and query-video contrastive learning; (vi) SSRN (Zhu et al. 2023) enriching anchor frames with additional consecutive frames; (vii) G2L (Li et al. 2023) measuring moment-query similarities using geodesic distance and quantifies cross-modal interactions with game-theoretic interactions; (viii) SOONet (Pan et al. 2023), an anchor-based framework that conducts grounding by pre-ranking, re-ranking, and regression; (ix) MESM (Liu et al. 2024), a fine-grained moment-query contrastive approach modeled for query word and video moment representations; (x) Contrastive-MSAT (Panta et al. 2024), applying moment-query contrastive loss supported by a momentum-based memory bank; (xi) UVCOM (Xiao et al. 2024), a moment-query contrastive approach for a unified video comprehension framework; (xii) SnAG (Mu, Mo, and Li 2024) achieving scalable grounding with cross-modal late fusion.
@@
In Figure 2, we observe that our model does not encounter degraded performance when the lengths of the target moments increase. Moreover, we visualize moment predictions of the recent method, i.e. SnAG (Mu, Mo, and Li 2024), and our model in Figure 1. Even though SnAG could precisely detect the shorter-length moment, it misses the moment of longer length, due to the degraded information issue. In contrast, our framework is able to localize both the short and long moments. We hypothesize that our contrastive framework can hold salient semantics for video moment representations to resolve the degraded signals in the grounding model, thus enhancing the grounding operation towards long video moments.

### v3

@@ Introduction
Temporal video grounding aims to localize moments of interest in an untrimmed video given a free-form textual description. It is a challenging multimodal task since it involves understanding temporal information in videos and reasoning about their connections to semantic information in texts. Recently, temporal grounding has drawn increasing attention (Mu et al. 2024; Jung et al. 2023; Xu et al. 2023; Pan et al. 2023), due to its wide range of applications such as surveillance (Zhang et al. 2016), robotics (Burgner-Kahrs et al. 2015), and autonomous driving (Claussmann et al. 2019).
@@
Previous methods (Zhang et al. 2020b; Soldan et al. 2021; Zhang et al. 2020a) for temporal grounding concentrate on grounding merely a few queries in short video snippets. However, recently the growing availability of long videos, e.g. on streaming platforms, and demands to query their rich content have necessitated productive grounding of large volumes of queries in long videos. Because of such short-to-long video paradigm shift, latest methods (Zhang et al. 2022; Mu et al. 2024) have utilized local self-attention to restrict attention within a local window, following the intuition that temporal context beyond a certain range is less helpful for moment localization.
@@
To capture moments at different temporal scales without enlarging the window size of the local self-attention, recent methods (Zhang et al. 2022; Mu et al. 2024) need to combine several Transformer blocks with downsampling between every two blocks, resulting in a feature pyramid of moment representations, as illustrated in Figure 1 (left). Unfortunately, due to such downsampling operation, when moment representations are propagated from lower levels of short-range (local) moments to higher levels of long-range (global) moments, information contained in representations of longer moments will gradually degrade (Guo et al. 2020; Yang et al. 2023). This could explain why performance of these methods tends to degrade as the duration of target moments increase, as shown in Figure 1 (right) and statistically shown with Intersection-over-Union (IoU) results in Figure 2, respectively.
@@
Figure 1: (Left) Illustration of feature pyramid to encode video moments of different lengths; (Right) An example where recent method SnAG (Mu et al. 2024) accurately localizes short video moment but fails on long moment.
@@
Figure 2: First and Second: IoU results with respect to target video moment length on Ego4D-NLQ (Grauman et al. 2022) of baseline SnAG (Mu et al. 2024) and our model. Third and Fourth: IoU results with respect to target video moment length on TACoS (Regneri et al. 2013) datasets of baseline SnAG (Mu et al. 2024) and our model.
@@
To enrich information in video moment representations, recent works (Panta et al. 2024; Xiao et al. 2024; Ji et al. 2024; Liu et al. 2024) have employed contrastive learning for temporal grounding. The intuition is to capture mutual information between video moments and textual query to preserve salient semantics in moment representations. These works mainly involve query-moment pairs in which queries relate to video moments of distinct videos, hence the learned semantics among moment representations would be independent from each other. However, such approach might not be suitable for the latest scalable video-centric approach (Zhang et al. 2022; Mu et al. 2024), in which multiple textual queries are related to one video. Therefore, if the grounding of two textual queries results in temporal overlapping, there might be a conflict in compact moment representations (An et al. 2023). Furthermore, focusing upon moment-query relations limits these works to the feature space of the final encoder layer, which could not effectively utilize all hidden representations across encoding layers. For multi-scale temporal grounding, such cross-scale representations should be fully used since they express semantics in video moments of various lengths.
@@
Temporal Grounding. Research works on temporal grounding can be categorized into two groups: two-stage and single-stage. In the two-stage group, methods generate temporal segments as proposals, then score the segments’ probabilities of being target moments and predict the refined boundary timestamps. Early approaches (Anne Hendricks et al. 2017; Gao et al. 2017) densely sample proposals leveraging sliding windows and score the proposals independently. Instead of independent sampling, Liu et al. (2021); Xiao et al. (2021a, b) subsequently condition proposal generation upon sentence queries and/or video context to avoid dense sampling. In contrast, Gao et al. (2021); Soldan et al. (2021); Wang et al. (2021a); Zhang et al. (2021) enumerate all segments and organize them into a 2D adjacency map for relation prediction. In the single-stage group, methods localize moments in a single shot without utilizing proposals, thus being more efficient than the two-stage group. Several works decode moment boundaries from a pooled representation (Li et al. 2022, 2021; Zhou et al. 2021) or learnable queries (Nguyen et al. 2023b; Lin et al. 2023; Woo et al. 2022). Most related to our work are models that denote moment candidates as points (Fang et al. 2023; Liu et al. 2022; Mu et al. 2024; Li et al. 2021). To capture temporal moments of various lengths, these methods designate a feature pyramid to produce moment representations of various scales, resulting in semantic gaps for longer-length moments.
Contrastive Learning. Notable improvements contributed by contrastive learning are made by contrastive loss applied to the final encoded outputs (Nguyen and Luu 2021; Wu et al. 2023; Nguyen et al. 2025; Hu et al. 2021; Nguyen et al. 2022; Wu et al. 2024; Nguyen et al. 2024b; Wang et al. 2021b; Nguyen et al. 2023a, 2024a). Wang et al. (2021b); Hu et al. (2021) employ a memory bank to maintain an extended set of positive and negative samples. Instead of utilizing outputs at the single final layer, contrastive learning with local and global representations across different layers has been widely studied (Hjelm et al. 2018; Zhang et al. 2020c; Bachman et al. 2019; Chaitanya et al. 2020). Zhang et al. (2020c) maximize the mutual information between representations of different local windows of a sentence and the representation of the global sentence. In addition, Bachman et al. (2019) propose an InfoNCE loss which optimizes over local and global features of two augmented views of an image, whereas Chaitanya et al. (2020) specifically concentrate on medical images.
@@
$\hskip-5.0pt\bar{Z}^{l}=\alpha^{l}\cdot\text{LocalMSA}\left(\text{LN}\left(Z^{l-1}\right)\right)+Z^{l-1},$ (2)
@@
Inspired by (Mu et al. 2024), we implement the downsampling operator $\downarrow$ as a strided depthwise 1D convolution. The downsampling operation engenders the multi-scale property of the encoder, generating representations for longer video moments.
@@
where $\mathcal{U}$ denotes a discrete uniform distribution, $\{y_{j^{\prime}}^{l}\}_{l=1}^{L}$ the set of target video moments in each layer $l$, and $N_{Q}$ the number of textual queries related to video $V$. We generate the target set $\mathcal{P}(l)$ via center sampling (Zhang et al. 2022; Mu et al. 2024), i.e. given any moment centered at $t$, any time step $c\in\left[t-\alpha\frac{T}{T^{l}},t+\alpha\frac{T}{T^{l}}\alpha\right]$ in layer $l$ is considered as a target. After sampling the query and target moments, we directly utilize the representations $\{\mathbf{z}_{j^{\prime}}^{l}\}_{l=1}^{L}$ of the target moments $\{y_{j^{\prime}}^{l}\}_{l=1}^{L}$ extracted by the aforementioned multi-scale video encoder.
@@
[…]
@@
We consider the following temporal grounding models as baselines: (i) VSL-Net (Zhang et al. 2020a) utilizing textual query to highlight regions potential to comprise the target moment; (ii) VLG-Net (Soldan et al. 2021) modeling temporal grounding as a graph matching problem; (iii) Moment-DETR (Lei et al. 2021), a Transformer encoder-decoder architecture that views temporal grounding as a set prediction problem; (iv) CONE (Hou et al. 2022) subsequently slicing a video input into windows, selects relevant windows, and ranks the selected windows to obtain target moments; (v) MMN (Wang et al. 2022), a Siamese-like network architecture that is trained with video-query and query-video contrastive learning; (vi) SSRN (Zhu et al. 2023) enriching anchor frames with additional consecutive frames; (vii) G2L (Li et al. 2023) measuring moment-query similarities using geodesic distance and quantifies cross-modal interactions with game-theoretic interactions; (viii) SOONet (Pan et al. 2023), an anchor-based framework that conducts grounding by pre-ranking, re-ranking, and regression; (ix) MESM (Liu et al. 2024), a fine-grained moment-query contrastive approach modeled for query word and video moment representations; (x) Contrastive-MSAT (Panta et al. 2024), applying moment-query contrastive loss supported by a momentum-based memory bank; (xi) UVCOM (Xiao et al. 2024), a moment-query contrastive approach for a unified video comprehension framework; (xii) SnAG (Mu et al. 2024) achieving scalable grounding with cross-modal late fusion.
@@
In Figure 2, we observe that our model does not encounter degraded performance when the lengths of the target moments increase. Moreover, we visualize moment predictions of the recent method, i.e. SnAG (Mu et al. 2024), and our model in Figure 1. Even though SnAG could precisely detect the shorter-length moment, it misses the moment of longer length, due to the degraded information issue. In contrast, our framework is able to localize both the short and long moments. We hypothesize that our contrastive framework can hold salient semantics for video moment representations to resolve the degraded signals in the grounding model, thus enhancing the grounding operation towards long video moments.